Physical Chemistry

Calculate the mass of a photon of sodium light having wavelength 5894 * 10^-10 and velocity 3 * 10^8m/s,h = plank’s constant.

Meghendra Agrawal
7 Years agoGrade

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
Dear Meghendra
 
"Wavelength=h/mv
so m=h/(wavelength*velocity)
=6.6*10^34/{[5894*10^(-10)]*3*10^8}
on solving we get
m=0.000373*10^(-32)

l = h / mc. l stands for wavelength 
So m = h / c l 
Hence m = 6.626*10^-34 / 3*10^8 * 589*10^-9 
So m = 3.75*10^-36 kg 
Compare this with atomic mass unit which is of the order of 10^-27 
Hence mass of photon is not considerable